OpenStudy (anonymous):

what is equal to 4^2*4^8

Nnesha (nnesha):

when we multiply same bases we should `add` exponents \[\huge\rm x^m \times x^n=x^{m+n}\]

OpenStudy (anonymous):

it is 65536+16=65552

Nnesha (nnesha):

oh-huh we don't even need calculator for this question you can solve it in few sec just add their exponent base would stay the same :=))